OpenStudy (anonymous):

I have the answer can someone check me please.. Find the derivative of f(x)=(5/square rootx)

OpenStudy (anonymous):

\[f(x)=\frac{ 5 }{ \sqrt{x} }\]

OpenStudy (phi):

write is like this \[ f(x) = 5 x^{-\frac{1}{2}} \] and use the "power rule"

OpenStudy (phi):

\[\frac{d}{dx} a x^n = a\ n\ x^{n-1} dx \]

OpenStudy (anonymous):

I got \[-\frac{ 5 }{ 2x ^{\frac{ 3 }{ 2 }} }\]

OpenStudy (phi):

\[ 5\cdot - \frac{1}{2} x^{- \frac{1}{2} - \frac{2}{2} } = -\frac{5}{2} x^{-\frac{3}{2}}\] which looks like your answer.

OpenStudy (anonymous):

ok so then that is the answer right?

OpenStudy (phi):

yes. x^(-3/2) can be written as 1/ x^(3/2)

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

Real quick questiondoes the negative go on the inside or outside of parenthesis -(5.....or (-5.......

OpenStudy (phi):

what parens?

OpenStudy (anonymous):

when typing the answer is it (-5/2x^(3/2))or -(5/2x(3/2))

OpenStudy (phi):

I would type (5/2) in parens so it is clear we have 5/2 the - can go in lots of places: (-5/2) or -(5/2) or (5/-2)

OpenStudy (anonymous):

ok thanks

OpenStudy (phi):

in your case where you have the x term in the denominator, I would write it -5/(2 x^(3/2)) with the parens around everything in the denominator

OpenStudy (anonymous):

ok thanks

OpenStudy (phi):

you could write it as \[ -\frac{5}{2 x^{\frac{3}{2}}} =\frac{-5}{2 x^{\frac{3}{2}}} =\frac{5}{-2 x^{\frac{3}{2}}}\]